Task Date Budgeted Amount Build forms April 1 $10,000 Pour foundation April 1 $50,000 May 1...

Task

Date

Budgeted Amount

Build forms

April 1

$10,000

Pour foundation

April 1

$50,000

May 1

$100,000

Frame walls

May 1

$30,000

June 1

$30,000

Remaining tasks

July 1 and beyond

$500,000

  1. SV as of May 1.
  2. Actual cost as of May 1 is $160,000. Calculate the cost variance (CV) as of May 1.
  3. Schedule performance index (SPI)
  4. Cost performance index (CPI)
  5. Estimate to complete (ETC), assuming that the previous cost variances will not affect future costs
  6. Estimate at completion (EAC)
Answer #1

As on May 1, the below ratios are calculated:

SV = Earned Value - Planned Value

SPI = Earned value/Planned value

Cost Performance Index = (Earned Value) / (Actual Cost)

Cost Variance = Budgeted cost - Actual cost

Planned value-

% completion as per Budgets = Budgeted cost as on May 1 / Total project cost = 190000/720000 = 0.263889 =~0.26

Budgeted cost * % completion as per Budgets = 190000*0.26 = 50,139

Earned value -

% completion as per Actuals = Actual cost as on May 1 / Total project cost = 160000/720000 = 0.22

Budgeted cost * % completion as per Actuals = 190000*0.22 = 42,222

SV = 42,222 - 50,139 = -7917 (behind schedule)

SPI = 42,222/50,139 = 0.84

CPI = 42,222/1,60,000 = 0.26